OpenStudy (anonymous):

The function below shows the number of car owners f(t), in thousand, in a city in different years t: f(t) = 0.25t2 - 1.5t + 2.5 The average rate of change of f(t) from t = 5 to t = 8 is ______ thousand per year

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yeah.. lol

OpenStudy (anonymous):

what can help you with

OpenStudy (anonymous):

the problem..lol

OpenStudy (anonymous):

are awarding medals

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

yeah...

jigglypuff314 (jigglypuff314):

plug in t=5 into the equation get number then plug in t=8 into equation and get another number then find slope (avg rate of change) which is second number minus first number divided by 8-5

OpenStudy (the_fizicx99):

Average rate of change is f(b) - f(a)/b-a

OpenStudy (the_fizicx99):

f(b) is 8 and f(a) is 5 Find the value of (5) and f(8) first

OpenStudy (anonymous):

f(t)=0.25t2−1.5t+2.5?

OpenStudy (anonymous):

have no idea

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

simplify 0.25(5)^2 - 1.5(5) + 2.5 then simplify 0.25(8)^2 - 1.5(8) + 2.5

OpenStudy (anonymous):

1.25,6.5

jigglypuff314 (jigglypuff314):

yep then do 6.5-1.5 --------- = ? 8 - 5

OpenStudy (the_fizicx99):

Remember to first do the exponents, 5^2 = 25 0.25*25 = 6.25 So f(5) = 6.25

OpenStudy (the_fizicx99):

Oh i missed the 1.5t >_<

OpenStudy (anonymous):

@jigglypuff314 5/3?

jigglypuff314 (jigglypuff314):

yeah, I think that's it :)

OpenStudy (anonymous):

your great thanks !

OpenStudy (the_fizicx99):

f(5) = 0.25(5)^2 - 1.5(5) + 2.5 f(5) = 0.25*25 - 7.5 + 2.5 f(5) = 6.25 - 7.5 + 2.5 f(5) = 1.25 Same for f(8), f(8) = 6.5 6.5 - 1.25 ---------- 8 - 5 5.25 ----- 3 So the average rate of change is 1.75 from 5 to 8 or 8 to 5

jigglypuff314 (jigglypuff314):

ah right, mistake on my part @desmonds @tHe_FiZiCx99 is right :)

OpenStudy (anonymous):

so it's 1.75?

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

I think so :3

OpenStudy (anonymous):

lol you think :x

OpenStudy (the_fizicx99):

What're that options :> if you have a doubt I'll double check it >_<

OpenStudy (anonymous):

lol ok i'll go with 1.75

OpenStudy (the_fizicx99):

o.0

OpenStudy (anonymous):

why the face lol?

OpenStudy (the_fizicx99):

Thought you had option >.>

OpenStudy (the_fizicx99):

options*